OpenStudy (anonymous):

Find S8 for the series -2+-10+-50+-250?

OpenStudy (anonymous):

PLEASE HELP!

OpenStudy (texaschic101):

are you looking for the 8th term or the sum of the terms up till the 8th term ?

OpenStudy (anonymous):

the 8th term of the sequence

OpenStudy (texaschic101):

Are you sure you are not looking for the sum of the first eight terms.....because that is what " s " means ? If you were looking for the 8th term, it would say " a8" of the series. So which one is it ? S8 or a8 ?

OpenStudy (anonymous):

could you possibly give me both answers?

OpenStudy (texaschic101):

yes...hold on...give me a moment

OpenStudy (anonymous):

k (:

OpenStudy (texaschic101):

The 8th term is : an = a1 * r^(n-1) n = 8 a1 = -2 r = 5 a8 = -2 * 5^(8 - 1) a8 = -2 * 5^7 a8 = - 2 * 78125 a8 = -156250 The sum is : sn = (a1(1 - r^n)/(1 - r) n = 8 a1 = -2 r = 5 s8 = (-2(1 - 5^8)/(1 - 5) s8 = (-2(1 - 390625)/-4 s8 = (-2(-390624)/-4 s8 = 781248/-4 s8 = -195312

OpenStudy (texaschic101):

I don't know which one you need, but here is both :)

OpenStudy (anonymous):

it was the sum! thanks! (:

OpenStudy (texaschic101):

no problem :)